# Defense Against Prompt Injection Attack by Leveraging Attack Techniques

### v3

@@ Abstract
With the advancement of technology, large language models (LLMs) have achieved remarkable performance across various natural language processing (NLP) tasks, powering LLM-integrated applications like Microsoft Copilot. However, as LLMs continue to evolve, new vulnerabilities, especially prompt injection attacks arise. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in data content, such as retrieved results. Recent attack methods leverage LLMs’ instruction-following abilities and their inabilities to distinguish instructions injected in the data content, and achieve a high attack success rate (ASR). When comparing the attack and defense methods, we interestingly find that they share similar design goals, of inducing the model to ignore unwanted instructions and instead to execute wanted instructions. Therefore, we raise an intuitive question: Could these attack techniques be utilized for defensive purposes? In this paper, we invert the intention of prompt injection methods to develop novel defense methods based on previous training-free attack methods, by repeating the attack process but with the original input instruction rather than the injected instruction. Our comprehensive experiments demonstrate that our defense techniques outperform existing defense approaches, achieving state-of-the-art results.
@@
Yulin Chen1thanks: Yulin Chen and Haoran Li contributed equally., Haoran Li211footnotemark: 1, Zihao Zheng3, Yangqiu Song2, Dekai Wu2, Bryan Hooi1 1National University of Singapore 2The Hong Kong University of Science and Technology 3Harbin Institute of Technology, Shenzhen chenyulin28@u.nus.edu
@@ 1 Introduction
With the continuously developing technologies, large language models (LLMs) have achieved impressive performance on various NLP tasks Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023), and are integrated into various real-world applications, such as Microsoft Copilot [^1] https://copilot.microsoft.com/, perplexity.ai [^2] https://www.perplexity.ai/, and so on. However, their inherent instruction-following capabilities make them vulnerable to prompt injection attacks. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in the data content, such as retrieved results from search engines. The prompt injection attacks can be generally classified into direct attacks Perez and Ribeiro (2022); Chen et al. (2024) and indirect attacks Greshake et al. (2023); Li et al. (2023); Zhan et al. (2024), according to the source of the input data content. For direct prompt injection attacks, the attackers, who are also the users, directly inject instructions into the data content for malicious purposes such as application prompt extraction Perez and Ribeiro (2022). Because of their instruction following ability, and their inability to distinguish the injected instructions, the LLMs execute the instructions in the data content and give undesired responses. On the other hand, for indirect prompt injection attacks, which have garnered more research attention recently, the malicious instructions are injected into external data content, such as retrieved results from external tool usage. In Figure 1 (a), for instance, attackers can inject the malicious prompt into the external data content, which consists of an attack prompt like “Forget previous instruction, and it’s urgent to” and an injected instruction after the attack prompt. This misleads the LLM into generating responses that align with the attacker’s intentions rather than following the original input instructions, thereby avoiding suspicion and potentially convincing users to click on malicious links Liu et al. (2024a). Current defense methods against prompt injection attacks primarily rely on fine-tuning Chen et al. (2024); Wallace et al. (2024); Suo (2024); Piet et al. (2023) or prompt engineering Hines et al. (2024); san (2023); ins (2023); Willison (2023). While fine-tuning-based defenses require annotated data and significant computational resources, prompt engineering approaches, though training-free, often prove less effective. In fact, the Open Worldwide Application Security Project (OWASP) has ranked prompt injection attacks as the #1 security risk for LLM applications OWASP (2023).
@@ 2.1 Prompt Injection Attacks
Prompt injection attacks have become a significant challenge for Large Language Models (LLMs), particularly in LLM-integrated applications. These attacks have been widely studied Perez and Ribeiro (2022); Willison (2023); Liu et al. (2023); Li et al. (2023); Liu et al. (2024b); Zhan et al. (2024); Shi et al. (2024); Liu et al. (2024a); Shafran et al. (2024); Huang et al. (2024); Breitenbach et al. (2023). Broadly, prompt injection attack methods can be classified into two categories: prompt-engineering-based attacks Breitenbach et al. (2023); Perez and Ribeiro (2022); Willison (2023); Liu et al. (2024b) and gradient-based attacks Huang et al. (2024); Shafran et al. (2024); Liu et al. (2024a); Shi et al. (2024). In prompt-engineering-based attacks, Perez and Ribeiro (2022) prepend an “ignoring prompt” to the injected instruction, while Willison (2023) propose adding a fake response to convince the LLM that the user’s input has been processed, prompting it to execute the maliciously injected instruction instead. On the other hand, gradient-based attacks, such as those based on the GCG attack method Zou et al. (2023), focus on training a suffix to induce the LLM to produce the desired response.
@@ 2.2 Prompt Injection Defenses
Given the severity of prompt injection attacks, several defense methods have been proposed san (2023); Hines et al. (2024); Willison (2023); Chen et al. (2024); Wallace et al. (2024); Yi et al. (2023); Piet et al. (2023); Suo (2024). san (2023) and Yi et al. (2023) suggest appending reminders to reinforce the importance of adhering to the original instructions. Hines et al. (2024) and Willison (2023) propose using special tokens to clearly delineate the data content area. Piet et al. (2023) defend against attacks by training models to perform specific tasks, rendering them incapable of following other potentially malicious instructions. Chen et al. (2024) and Wallace et al. (2024) advocate fine-tuning LLMs with instruction-following datasets, granting privileged status to authorized instructions. Lastly, Suo (2024) introduce a method of signing instructions with special tokens, ensuring that LLMs only follow those that are properly signed.
@@ 3.4 Fake Completion Attack
However, this example represents a relatively weaker attack, as it assumes that the attacker does not have knowledge of the full conversation template. For instance, in the case of Figure 13, the attacker uses ‘###instruction:’ as the instruction identifier, whereas the actual identifier is ‘<Instruction>’. If the attacker has access to the entire conversation template, they can fabricate a more convincing assistant response, as illustrated in Figure 14, making this type of attack much harder to defend against.
@@ 4.2.3 Fake Completion Defense
Another attack method, the fake completion attack, similarly misleads the model into ignoring the original instruction. In this attack, a fake response (attack prompt) such as ‘### Response: OK’ tricks the LLM into believing that the original instruction has been completed. For our defense, we mimic this approach by repeating the fake response as our shield prompt. As shown in Figure 2 (c), we first fabricate a response to the last instruction. To keep it simple, we use the same response as the attacker’s fake response—‘OK’. This defensive response can evolve alongside improvements in attack techniques. We then append the original input instruction. Upon reading the fabricated response, the LLM will assume the injected instructions have already been executed and will only follow the appended original input instruction. Since the fake completion defense simulates a multi-turn conversation, it could be possible to design the defense within a real multi-turn conversation structure.
@@ 5.1 Experimental Settings
We evaluate our defense methods against both direct and indirect prompt injection attacks. For direct injection attacks, we follow the method of Chen et al. (2024), applying attacks to 208 samples from AlpacaFarm Dubois et al. (2024) and comparing the effectiveness of our defense methods with baseline approaches. For indirect prompt injection attacks, we use the QA dataset filtered by Li et al. (2023), where malicious instructions are injected into retrieved data content for evaluation, and this dataset contains 2000 samples.
@@
We select popular and strong open-source LLMs as victim models for our experiments. Specifically, we choose Llama3.1-8b-Instruct Dubey et al. (2024), Qwen2-7b-Instruct Yang et al. (2024), and Llama3-8b-Instruct AI@Meta (2024). Throughout the experiments, unless otherwise specified, “Llama3”, “Llama3.1” and “Qwen2” refer to Llama3-8b-Instruct, Llama3.1-8b-Instruct, and Qwen2-7b-Instruct, respectively.
@@
In our experimental setup, we assume that for our methods, only the utilized attack method is known during defense, and all other attack methods remain unknown. This setup challenges the generalization ability of our methods. For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detecte if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023) and the sentiment analysis dataset SST2 Socher et al. (2013), which are not attacked and include the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@ 5.2.1 Attack Methods
As discussed in Section 3, we select the following attack methods for evaluation: Naive attack (abbreviated as “Naive”), Ignore attack (“Ignore”), Escape-Deletion attack (“Escape”), Fake completion attack (“Fakecom”), and Fake completion attack with template (“Fakecom-t”). Additionally, we include a Combined attack Liu et al. (2024b), which combines the Ignore attack, Fake completion attack, and Escape-Separation attack, referred as “Combined.” An example is shown in Figure 15.
@@ 5.3.2 Defense against Indirect Attack
In addition to evaluating defense against direct prompt injection attacks, we also assess its effectiveness against indirect attacks. The key difference between direct and indirect prompt injection attacks is that, in the case of indirect attacks, the input data is retrieved from external tools, such as search engines, and users are often unaware of the attack. To evaluate indirect prompt injection attacks, we use the filtered QA dataset from Li et al. (2023). Table 2 shows the results of our defense methods compared to the baselines in the indirect scenario. Our methods continue to outperform the baselines by a significant margin. When comparing both direct and indirect prompt injection attacks, it appears that indirect attacks are easier to defend against. Furthermore, Qwen2 remains the most susceptible model to attacks compared to the other two models.
@@ 5.3.3 Model Utility
A key evaluation metric for defense methods is their potential impact on the model’s utility. To assess the impact of our method, we use the filtered QA dataset from Li et al. (2023). For simplicity, we do not introduce any attacks into the retrieved data content, and we only verify whether the correct (golden) answer appears in the model’s response, with different defense methods. Table 3 presents the utility performance of various defense strategies. Notably, most defense strategies do not significantly affect the model’s utility. Moreover, our proposed defense methods can even improve the performance in some scenarios. Additionally, to further validate the robustness of our results, we conduct experiments on the sentiment analysis task using the SST2 dataset Socher et al. (2013), with results shown in Table 10. The results demonstrate that our methods cause minimal degradation to the model’s overall performance.
@@ 5.4 Ablation Study
Beyond prompt-engineering-based attacks, we also evaluate the effectiveness of our defense methods against gradient-based attacks. Specifically, we perform direct prompt injection attacks using the GCG method Zou et al. (2023) and the AutoDAN method Zhu et al. (2023) with Llama3. Table 4 presents the defense results. Our first observation is that compared to baseline methods, our defense strategies more effectively mitigate these attacks. Notably, the “Fakecom-t” method proves to be the most effective, reducing the ASR to around 10% and demonstrating strong transferability across different attack types.
@@
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the “Naive attack” and the “Ignore attack” respectively for evaluation. Table 6 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because “Ignore attack” is a part of “Combined attack”, “StruQ-Ignore” can defend against “Combined attack” successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
@@
Table 6: Defense performance of our methods and the fine-tuning method StruQ. “StruQ-Naive” means StruQ incorporates the “Naive attack” for fine-tuning. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
@@
We also examine the impact of deleting data content. For more details, please refer to Appendix A.3.
@@ 5.5 Case Study
Figure 16 provides two examples of responses without defense against the “Ignore attack” and “Fakecom attack”, both generated by Llama3. Additionally, we include the response generated by the model with defense method based on “Fakecom-t”, as it is the most effective defense approach. From the examples, we observe that the “Ignore attack” does not consistently persuade the LLM to ignore the original input instruction and the model may end up executing both instructions. Although in this instance, the “Fakecom attack” successfully misleads the LLM to execute the injected instruction directly, this strategy does not always work, and there are cases where the model executes both instructions, explaining the failures of the defense methods based on these attacks. In the case of the defense method based on “Fakecom-t”, we can observe that the defense method successfully enables the LLM to bypass the injected instruction. What’s more, the response still remains relevant to the original task, suggesting the defense method has little damage on the utility of the model.

### v4

@@ Abstract
With the advancement of technology, large language models (LLMs) have achieved remarkable performance across various natural language processing (NLP) tasks, powering LLM-integrated applications like Microsoft Copilot. However, as LLMs continue to evolve, new vulnerabilities, especially prompt injection attacks arise. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in data content, such as retrieved results. Recent attack methods leverage LLMs’ instruction-following abilities and their inabilities to distinguish instructions injected in the data content, and achieve a high attack success rate (ASR). When comparing the attack and defense methods, we interestingly find that they share similar design goals, of inducing the model to ignore unwanted instructions and instead to execute wanted instructions. Therefore, we raise an intuitive question: Could these attack techniques be utilized for defensive purposes? In this paper, we invert the intention of prompt injection methods to develop novel defense methods based on previous training-free attack methods, by repeating the attack process but with the original input instruction rather than the injected instruction. Our comprehensive experiments demonstrate that our defense techniques outperform existing defense approaches, achieving state-of-the-art results. [^1] Code is publicly available at https://github.com/LukeChen-go/pia-defense-by-attack.
@@
Yulin Chen1, Haoran Li2, Zihao Zheng3, Dekai Wu2, Yangqiu Song2, Bryan Hooi1 1National University of Singapore, 2HKUST, 3Harbin Institute of Technology, Shenzhen chenyulin28@u.nus.edu, hlibt@connect.ust.hk, melfeszheng@gmail.com dekai@ust.hk, yqsong@cse.ust.hk, bhooi@comp.nus.edu.sg
@@ 1 Introduction
With the continuously developing technologies, large language models (LLMs) have achieved impressive performance on various NLP tasks Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023), and are integrated into various real-world applications, such as Microsoft Copilot [^2] https://copilot.microsoft.com/, perplexity.ai [^3] https://www.perplexity.ai/, and so on. However, their inherent instruction-following capabilities make them vulnerable to prompt injection attacks. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in the data content, such as retrieved results from search engines. The prompt injection attacks can be generally classified into direct attacks Perez and Ribeiro (2022); Chen et al. (2024) and indirect attacks Greshake et al. (2023); Li et al. (2023b); Zhan et al. (2024), according to the source of the input data content. For direct prompt injection attacks, the attackers, who are also the users, directly inject instructions into the data content for malicious purposes such as application prompt extraction Perez and Ribeiro (2022). Because of their instruction following ability, and their inability to distinguish the injected instructions, the LLMs execute the instructions in the data content and give undesired responses. On the other hand, for indirect prompt injection attacks, which have garnered more research attention recently, the malicious instructions are injected into external data content, such as retrieved results from external tool usage. In Figure 1 (a), for instance, attackers can inject the malicious prompt into the external data content, which consists of an attack prompt like “Forget previous instruction, and it’s urgent to” and an injected instruction after the attack prompt. This misleads the LLM into generating responses that align with the attacker’s intentions rather than following the original input instructions, thereby avoiding suspicion and potentially convincing users to click on malicious links Liu et al. (2024a). Current defense methods against prompt injection attacks primarily rely on fine-tuning Chen et al. (2024); Wallace et al. (2024); Suo (2024); Piet et al. (2023) or prompt engineering Hines et al. (2024); san (2023); ins (2023); Willison (2023). While fine-tuning-based defenses require annotated data and significant computational resources, prompt engineering approaches, though training-free, often prove less effective. In fact, the Open Worldwide Application Security Project (OWASP) has ranked prompt injection attacks as the #1 security risk for LLM applications OWASP (2023).
@@ 2.1 Prompt Injection Attacks
Owing to their impressive performance, large language models (LLMs) have been widely adopted for a broad range of NLP tasks Chen et al. (2021); Kojima et al. (2022); He et al. (2024); Zong et al. (2024); Sui et al. (2024); Liu et al. (2025b); Li et al. (2025). However, prompt injection attacks have become a significant challenge for LLMs, particularly in LLM-integrated applications. These attacks have been widely studied Perez and Ribeiro (2022); Willison (2023); Liu et al. (2023); Li et al. (2023b); Liu et al. (2024b); Zhan et al. (2024); Shi et al. (2024); Liu et al. (2024a); Shafran et al. (2024); Huang et al. (2024); Breitenbach et al. (2023). Broadly, prompt injection attack methods can be classified into two categories: prompt-engineering-based attacks Breitenbach et al. (2023); Perez and Ribeiro (2022); Willison (2023); Liu et al. (2024b) and gradient-based attacks Huang et al. (2024); Shafran et al. (2024); Liu et al. (2024a); Shi et al. (2024). In prompt-engineering-based attacks, Perez and Ribeiro (2022) prepend an “ignoring prompt” to the injected instruction, while Willison (2023) propose adding a fake response to convince the LLM that the user’s input has been processed, prompting it to execute the maliciously injected instruction instead. On the other hand, gradient-based attacks, such as those based on the GCG attack method Zou et al. (2023), focus on training a suffix to induce the LLM to produce the desired response.
@@ 2.2 Prompt Injection Defenses
Given the severity of prompt injection attacks, several defense methods have been proposed san (2023); Hines et al. (2024); Willison (2023); Chen et al. (2024); Wallace et al. (2024); Zhan et al. (2024); Piet et al. (2023); Suo (2024); Li et al. (2023a); Liu et al. (2025a). san (2023) and Yi et al. (2023) suggest appending reminders to reinforce the importance of adhering to the original instructions. Hines et al. (2024) and Willison (2023) propose using special tokens to clearly delineate the data content area. Piet et al. (2023) defend against attacks by training models to perform specific tasks, rendering them incapable of following other potentially malicious instructions. Chen et al. (2024) and Wallace et al. (2024) advocate fine-tuning LLMs with instruction-following datasets, granting privileged status to authorized instructions. Lastly, Suo (2024) introduce a method of signing instructions with special tokens, ensuring that LLMs only follow those that are properly signed.
@@ 3.4 Fake Completion Attack
However, this example represents a relatively weaker attack, as it assumes that the attacker does not have knowledge of the full conversation template. For instance, in the case of Figure 13, the attacker uses “###instruction:” as the instruction identifier, whereas the actual identifier is “<Instruction>.” If the attacker has access to the entire conversation template, they can fabricate a more convincing assistant response, as illustrated in Figure 14, making this type of attack much harder to defend against.
@@ 4.2.3 Fake Completion Defense
Another attack method, the fake completion attack, similarly misleads the model into ignoring the original instruction. In this attack, a fake response (attack prompt) such as “### Response: OK” tricks the LLM into believing that the original instruction has been completed. For our defense, we mimic this approach by repeating the fake response as our shield prompt. As shown in Figure 2 (c), we first fabricate a response to the last instruction. To keep it simple, we use the same response as the attacker’s fake response—“OK.” This defensive response can evolve alongside improvements in attack techniques. We then append the original input instruction. Upon reading the fabricated response, the LLM will assume the injected instructions have already been executed and will only follow the appended original input instruction. Since the fake completion defense simulates a multi-turn conversation, it could be possible to design the defense within a real multi-turn conversation structure.
@@ 5.1 Experimental Settings
We evaluate our defense methods against both direct and indirect prompt injection attacks. For direct injection attacks, we follow the method of Chen et al. (2024), applying attacks to 208 samples from AlpacaFarm Dubois et al. (2024) and comparing the effectiveness of our defense methods with baseline approaches. For indirect prompt injection attacks, we use the QA dataset filtered by Li et al. (2023b), where malicious instructions are injected into retrieved data content for evaluation, and this dataset contains 2000 samples.
@@
We select popular and strong open-source LLMs as victim models for our experiments. Specifically, we choose Llama3.1-8b-Instruct Dubey et al. (2024), Qwen2-7b-Instruct Yang et al. (2024), and Llama3-8b-Instruct AI@Meta (2024). Throughout the experiments, unless otherwise specified, “Llama3,” “Llama3.1,” and “Qwen2” refer to Llama3-8b-Instruct, Llama3.1-8b-Instruct, and Qwen2-7b-Instruct, respectively.
@@
In our experimental setup, we assume that for our methods, only the utilized attack method is known during defense, and all other attack methods remain unknown. This setup challenges the generalization ability of our methods. For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detect if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023b) and the sentiment analysis dataset SST2 Socher et al. (2013), which are not attacked and include the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
@@ 5.2.1 Attack Methods
As discussed in Section 3, we select the following attack methods for evaluation: Naive attack (abbreviated as “Naive”), Ignore attack (“Ignore”), Escape characters attack (“Escape”), Fake completion attack (“Fakecom”), and Fake completion attack with template (“Fakecom-t”). Additionally, we include a Combined attack Liu et al. (2024b), which combines the Ignore attack, Fake completion attack, and Escape characters attack, referred as “Combined.” An example is shown in Figure 15.
@@ 5.3.2 Defense against Indirect Attack
In addition to evaluating defense against direct prompt injection attacks, we also assess its effectiveness against indirect attacks. The key difference between direct and indirect prompt injection attacks is that, in the case of indirect attacks, the input data is retrieved from external tools, such as search engines, and users are often unaware of the attack. To evaluate indirect prompt injection attacks, we use the filtered QA dataset from Li et al. (2023b). Table 2 shows the results of our defense methods compared to the baselines in the indirect scenario. Our methods continue to outperform the baselines by a significant margin. When comparing both direct and indirect prompt injection attacks, it appears that indirect attacks are easier to defend against. Furthermore, Qwen2 remains the most susceptible model to attacks compared to the other two models.
@@ 5.3.3 Model Utility
A key evaluation metric for defense methods is their potential impact on the model’s utility. To assess the impact of our method, we use the filtered QA dataset from Li et al. (2023b). For simplicity, we do not introduce any attacks into the retrieved data content, and we only verify whether the correct (golden) answer appears in the model’s response, with different defense methods. Table 3 presents the utility performance of various defense strategies. Notably, most defense strategies do not significantly affect the model’s utility. Moreover, our proposed defense methods can even improve the performance in some scenarios. Additionally, to further validate the robustness of our results, we conduct experiments on the sentiment analysis task using the SST2 dataset Socher et al. (2013), with results shown in Table 11. The results demonstrate that our methods cause minimal degradation to the model’s overall performance.
@@
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21
@@ 5.4 Ablation Study
Beyond prompt-engineering-based attacks, we also evaluate the effectiveness of our defense methods against gradient-based attacks. Specifically, we perform direct prompt injection attacks using the GCG method Zou et al. (2023) and the AutoDAN method Zhu et al. (2023) with Llama3. Table 5 presents the defense results. Our first observation is that compared to baseline methods, our defense strategies more effectively mitigate these attacks. Notably, our defense method based on “Fakecom-t” proves to be the most effective, reducing the ASR to around 10% and demonstrating strong transferability across different attack types.
@@
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods in indirect scenario with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the “Naive attack” and the “Ignore attack” respectively for evaluation. Table 7 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because “Ignore attack” is a part of “Combined attack,” “StruQ-Ignore” can defend against “Combined attack” successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
@@
Table 7: Defense performance of our method and the fine-tuning method StruQ. “StruQ-Naive” means StruQ incorporates the “Naive attack” for fine-tuning. The evaluation metric is ASR. Bold indicates the best performance. All results are reported in %.
@@ 5.5 Case Study
Figure 16 provides two examples of responses without defense against the “Ignore attack” and “Fakecom attack,” both generated by Llama3. Additionally, we include the response generated by the model with defense method based on “Fakecom-t,” as it is the most effective defense approach. From the examples, we observe that the “Ignore attack” does not consistently persuade the LLM to ignore the original input instruction and the model may end up executing both instructions. Although in this instance, the “Fakecom attack” successfully misleads the LLM to execute the injected instruction directly, this strategy does not always work, and there are cases where the model executes both instructions, explaining the failures of the defense methods based on these attacks. In the case of the defense method based on “Fakecom-t,” we can observe that the defense method successfully enables the LLM to bypass the injected instruction. What’s more, the response still remains relevant to the original task, suggesting the defense method has little damage on the utility of the model.